OpenStudy (anonymous):

Find the equation of the line containing points: -2,2 and 2,-2 arrange the answer in y=mx+b format where b is the constant!

OpenStudy (anonymous):

You can work out the gradient (m) from the points given. So m=(y2-y1)/(x2-x1)=(-2-2)/(2-(-2))=-1.